OpenStudy (anonymous):

how do you solve 3^(x – 4) = 7^(x + 9)

OpenStudy (hba):

use the log function

OpenStudy (anonymous):

can you show me how you would set up this problem

OpenStudy (anonymous):

okay \[\ln{3^(x-4)}\]=\[\ln{7^(x+9)}\] ln7*(x+9)=ln3*(x-4)

OpenStudy (anonymous):

so calculate ln3&ln7 by calculator

OpenStudy (anonymous):

r u agree